A restaurant chain has 4 locations around a city. Management wants to know if customers are spending roughly the same amount of time during the peak hours per table. Analyst collected data for 5 random tables at each loaction. That data is presented in the table below. For example, 47 means that a randomly selected table was occupied for 47 minutes to serve one group of patrons. Restarant Time Spent Per Table (During Peak Hours) Data Location A 52 75 67 58 59 Location B 55 77 38 81 95 Location C 47 82 63 46 74 Location D 77 73 85 56 91.

Hypothesis 2 :The average time spent per table in Location A is more than the average time spent per table in Location B.
Answer #1

The average time spent per table in location A

= ( 52+75+67+58+59 ) ÷ 5

=311÷5 = 62.2

The average time spent per table in location B

= ( 55+77+38+81+95 ) ÷ 5

= 346 ÷ 5 = 69.2

Thus, the hypothesis 2 is rejected. The average time spent per table in location A is less than the average time spent per table in location B.